# Synergy over Discrepancy: A Partition-Based Approach to Multi-Domain LLM Fine-Tuning

### v2

@@ 1 Introduction
Large language models (LLMs) have propelled natural language processing (NLP) to unprecedented capabilities (Kumar, 2024; Karanikolas et al., 2023; Hu et al., 2025; Zhang et al., 2025), owing largely to their extensive pretraining on massive, diverse textual corpora (Wu et al., 2022; Li et al., 2024a; Chen et al., 2025). Fine-tuning these pretrained models for a specific downstream domain has been widely explored and proven highly effective; notable approaches include adapter-based modules (Houlsby et al., 2019; Zhang et al., 2024a), parameter-efficient fine-tuning via low-rank updates (Hu et al., 2021), and instruction-based fine-tuning methods (Cao et al., 2024). These methods have also accelerated the application of LLMs in many scenarios, such as healthcare(Tong et al., 2025; Jiaqi Liu, 2025), transportation(Yao et al., 2023; Wang et al., 2025), and robotics(Xiao et al., 2025a; Li et al., 2025).
@@
However, while these methods excel in adapting to a single domain, practical scenarios frequently require simultaneous adaptation to multiple distinct domains-a scenario far less studied and substantially more challenging. Consider, for instance, a scenario where a single pretrained model must be simultaneously adapted to clinical texts (Thirunavukarasu et al., 2023), social media posts (Yang et al., 2024a; Jiang et al., 2024), and legal documents (Seabra et al., 2024). Naive approaches such as jointly fine-tuning the model across all domains or independently fine-tuning separate models per domain often yield suboptimal results(Xiao et al., 2025b): domain-specific features may negatively interfere, causing one domain to overshadow others or impair overall generalization capabilities (Lu et al., 2024; Zheng et al., 2024; Van Veen et al., 2023; Tan et al., 2025). This motivates our central research question: how can we effectively and efficiently fine-tune a single LLM across multiple heterogeneous domains, exploiting inter-domain synergies while mitigating negative interference?
@@
A natural strategy to mitigate these challenges is to add adapter modules specialized for each domain (Houlsby et al., 2019; Zhang et al., 2024a). While adapters reduce the need for full fine-tuning, they do not fully address the complexity of managing multiple source domains with potentially large discrepancies. In such cases, the learned model has to juggle both shared features (common linguistic properties across domains) and domain-specific features (rare words, styles, or content). Existing approaches often handle these competing demands by imposing either domain adversarial objectives (Ganin and Lempitsky, 2015), distribution alignment (Peng et al., 2019), or low-rank parameter updates (Hu et al., 2021). However, these techniques may still fail to exploit synergistic relationships-where certain domains are complementary and can reinforce each other’s accuracy-and do not fully capture how best to “partition” domains to avoid negative interference.
@@ 2.1 LLM Fine-Tuning
The success of LLMs such as GPT-3 (Brown et al., 2020), LLaMA (Touvron et al., 2023), and Falcon (Almazrouei et al., 2023) has underscored the importance of efficient fine-tuning. Fully updating model parameters incurs high computational costs (Zhang et al., 2024a; Hu et al., 2021), prompting parameter-efficient methods that modify only a small subset of parameters. Examples include adapter modules (Houlsby et al., 2019), low-rank projections (LoRA) (Hu et al., 2021), and tightly integrated adapters for LLaMA (Zhang et al., 2024a). However, most methods assume a single domain; adapting LLMs efficiently to multiple domains remains challenging.
@@ 2.2 Multi-Domain Data Learning
Real-world data often originate from disparate sources with distinct distributions and vocabularies (Zhang et al., 2024b; Ganin and Lempitsky, 2015; Sener and Koltun, 2018; Xiao, 2024; Xiao and Liu, 2025). Traditional multi-domain methods align representations through adversarial training (Pei et al., 2018; Ganin and Lempitsky, 2015), moment matching (Peng et al., 2019), or multi-task objectives (Royer et al., 2024), yet typically neglect inter-domain synergies. Adapting LLMs further complicates this via memory overhead, forgetting pretrained knowledge, and cross-domain contamination. Adapter-based solutions partially address these concerns (Houlsby et al., 2019) but rarely exploit domain partitioning to maximize synergy. Our partition-based multi-stage approach systematically clusters domains to leverage synergy and minimize discrepancy, providing theoretical guarantees.
@@ 2.3 LLM Data Selection
The efficacy of supervised fine-tuning (SFT) heavily depends on the quality and composition of training data(Yao et al., 2024). Recent advances have introduced diverse metrics for data selection: instance-level criteria like perplexity (Cao et al., 2024), reward scores (Gou and Nguyen, 2024), and loss disparities (Li et al., 2024b), trajectory-based clustering via small proxy models (Yang et al., 2024b), as well as token-level selection methods (Lin et al., 2024). Existing works largely overlook one critical factor: domain interactions, where diversity metrics operate at instance/token levels (Pang et al., 2024) without modeling cross-domain compatibility. Our approach addresses this gap by systematically clustering domains for joint tuning.
@@ 5 Experiments
LLaMA-Adapter (Zhang et al., 2024a) 41.5 64.3 89.1 86.7 42.6 65.9 89.7 87.5 43.1 67.3 90.3 88.1
@@
We compare PMS-FTP against the following baselines: 1) Base Methods: Full Fine-Tuning (FULL), Fixed Backbone (FIXED); 2) Domain Adaptation: Multi-Domain Adversarial Network (MDAN) (Pei et al., 2018), Moment Matching (M3SDA) (Peng et al., 2019), Bayesian Gaussian Mixture (GMDI) (Ling et al., 2024); 3) Single-Domain LLM Fine-tuning: LoRA (Hu et al., 2021), Adapter (Houlsby et al., 2019), LLaMA-Adapter (Zhang et al., 2024a), Q-LoRA (Dettmers et al., 2023), Tag-LLM (Shen et al., 2024); 4) Data Selection: INSTRUCTMINING (IT) (Cao et al., 2024), S2L (Yang et al., 2024b). Please refer to A.2 for more details on the baselines.
@@ 5.1 Experimental Results
Memory Usage.
Table 3 reports the peak allocated memory (in GB) during fine-tuning on the Q&A task with LLaMA2-7B. We measure memory usage using a single NVIDIA A100 GPU. The FULL method requires the largest memory footprint due to updating all model parameters. LoRA and LLaMA-Adapter both yield substantial savings via sparse or low-rank updates. Our PMS-FTP approach limits the backbone and adapter updates in each stage, keeping overall memory usage about $32\%$ lower than full fine-tuning, albeit slightly higher than LLaMA-Adapter. Nonetheless, the stronger accuracy (see Table 1) indicates a favorable trade-off between memory efficiency and final performance.
@@ 6 Conclusion
In future work, we aim to adapt this stage-wise procedure to a continual learning setting, where new domains arrive sequentially, thereby offering a more flexible lifelong learning framework for large language models.
@@ 10. Guidelines:
    • The authors are encouraged to create a separate "Limitations" section in their paper.
@@ 35. Guidelines:
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
@@ 60. Guidelines:
    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
@@ A.2 Baselines
    • LLaMA-Adapter [Zhang et al., 2024a]: zero-init residual adapters.
@@ A.3 Implementation Details
where $V_{i}$ is the vocabulary set of $\mathcal{D}_{i}$, $\operatorname{Jacc}(V_{i},V_{j})=|V_{i}\cap V_{j}|/|V_{i}\cup V_{j}|$, and $\mu_{i}$ is the mean Sentence-BERT embedding of domain ii. Both terms are in $[0,1]$; the average is therefore in $[0,1]$.
@@ B.7 Inference footprint after LoRA merging
After each stage we merge the finished LoRA into the frozen backbone, so only one 4-bit adapter is carried at inference. Table 15 shows equal-or-lower memory than a single-adapter baseline. Together with the accuracy gains in Table 1, merging achieves a strictly better accuracy–memory trade-off than training/keeping multiple adapters.
@@ B.10 Incremental addition of a new domain
After training on the original four domains, we add “Legal-QA” as a new stage and freeze prior adapters. Table 19 shows negligible forgetting on old domains and a gain over single-domain LoRA on the new domain. Disjoint, frozen adapters make PMS-FTP naturally amenable to one-shot domain extension without replay.
Table 19: One-shot incremental stage (add “Legal-QA”).

### v3

@@ 1 Introduction
Large language models (LLMs) have propelled natural language processing (NLP) to unprecedented capabilities (Kumar, 2024; Karanikolas et al., 2023; Hu et al., 2025; Zhang et al., 2025b; Lin et al., 2025), owing largely to their extensive pretraining on massive, diverse textual corpora (Wu et al., 2022; Li et al., 2024a; Chen et al., 2025b; Zhang et al., 2024a; Sun et al., 2025b). Fine-tuning these pretrained models for a specific downstream domain has been widely explored and proven highly effective; notable approaches include adapter-based modules (Houlsby et al., 2019; Zhang et al., 2024b), parameter-efficient fine-tuning via low-rank updates (Hu et al., 2021), and instruction-based fine-tuning methods (Cao et al., 2024). These methods have also accelerated the application of LLMs in many scenarios, such as healthcare(Tong et al., 2025; Liu et al., 2025; Wang et al., 2025), transportation(Yao et al., 2023; Lu et al., 2025b; Zeng et al., 2025b, a), robotics(Xiao et al., 2025a; Yan et al., 2025; Zhang et al., 2025a), and other applications (Sun et al., 2025a; Zhang et al., 2025c).
@@
However, while these methods excel in adapting to a single domain, practical scenarios frequently require simultaneous adaptation to multiple distinct domains-a scenario far less studied and substantially more challenging (Lu et al., 2025a; Chen et al., 2025a). Due to constraints on energy or computing power (He, 2025), we must consider how to enhance the model’s ability to adapt to different domains. Consider, for instance, a scenario where a single pretrained model must be simultaneously adapted to clinical texts (Thirunavukarasu et al., 2023), social media posts (Yang et al., 2024a; Jiang et al., 2024, 2025), and legal documents (Seabra et al., 2024). Naive approaches such as jointly fine-tuning the model across all domains or independently fine-tuning separate models per domain often yield suboptimal results(Xiao et al., 2025b; Zhang et al., 2023; Wang and Zhang, 2024; Wen et al., 2023): domain-specific features may negatively interfere, causing one domain to overshadow others or impair overall generalization capabilities (Lu et al., 2024a; Zheng et al., 2024; Van Veen et al., 2023; Tan et al., 2025). This motivates our central research question: how can we effectively and efficiently fine-tune a single LLM across multiple heterogeneous domains, exploiting inter-domain synergies while mitigating negative interference?
@@
A natural strategy to mitigate these challenges is to add adapter modules specialized for each domain (Houlsby et al., 2019; Zhang et al., 2024b; Tao et al., 2023). While adapters reduce the need for full fine-tuning, they do not fully address the complexity of managing multiple source domains with potentially large discrepancies. In such cases, the learned model has to juggle both shared features (common linguistic properties across domains) and domain-specific features (rare words, styles, or content) (Lu et al., 2022). Existing approaches often handle these competing demands by imposing either domain adversarial objectives (Ganin and Lempitsky, 2015), distribution alignment (Peng et al., 2019), or low-rank parameter updates (Hu et al., 2021). However, these techniques may still fail to exploit synergistic relationships-where certain domains are complementary and can reinforce each other’s accuracy-and do not fully capture how best to “partition” domains to avoid negative interference.
@@ 2.1 LLM Fine-Tuning
The success of LLMs such as GPT-3 (Brown et al., 2020), LLaMA (Touvron et al., 2023), and Falcon (Almazrouei et al., 2023) has underscored the importance of efficient fine-tuning. Fully updating model parameters incurs high computational costs (Zhang et al., 2024b; Hu et al., 2021), prompting parameter-efficient methods that modify only a small subset of parameters. Examples include adapter modules (Houlsby et al., 2019), low-rank projections (LoRA) (Hu et al., 2021), and tightly integrated adapters for LLaMA (Zhang et al., 2024b). However, most methods assume a single domain; adapting LLMs efficiently to multiple domains remains challenging. Another related field is continual learning (Xu et al., 2025), but it assumes a sequential arrival of tasks, which differs from the setting in this paper.
@@ 2.2 Multi-Domain Data Learning
Real-world data often originate from disparate sources with distinct distributions and vocabularies (Ganin and Lempitsky, 2015; Sener and Koltun, 2018; Li et al., 2023; Zhang et al., 2024c; Chen et al., 2024; Xiao and Liu, 2025; Ke et al., 2025). Traditional multi-domain methods align representations through adversarial training (Pei et al., 2018; Ganin and Lempitsky, 2015), moment matching (Peng et al., 2019), or multi-task objectives (Royer et al., 2024; Shen and Zhang, 2025), yet typically neglect inter-domain synergies. Adapting LLMs further complicates this via memory overhead, forgetting pretrained knowledge, and cross-domain contamination. Adapter-based solutions partially address these concerns (Houlsby et al., 2019) but rarely exploit domain partitioning to maximize synergy. Our partition-based multi-stage approach systematically clusters domains to leverage synergy and minimize discrepancy, providing theoretical guarantees.
@@ 2.3 LLM Data Selection
The efficacy of supervised fine-tuning (SFT) heavily depends on the quality and composition of training data(Yao et al., 2024). Recent advances have introduced diverse metrics for data selection: instance-level criteria like perplexity (Cao et al., 2024), reward scores (Gou and Nguyen, 2024), and loss disparities (Li et al., 2024b), trajectory-based clustering via small proxy models (Yang et al., 2024b), as well as token-level selection methods (Lin et al., 2024). Existing works largely overlook one critical factor: domain interactions, where diversity metrics operate at instance/token levels (Pang et al., 2024) without modeling cross-domain compatibility (Li et al., 2025b). Our approach addresses this gap by systematically clustering domains for joint tuning.
@@ 5 Experiments
LLaMA-Adapter (Zhang et al., 2024b) 41.5 64.3 89.1 86.7 42.6 65.9 89.7 87.5 43.1 67.3 90.3 88.1
@@
We compare PMS-FTP against the following baselines: 1) Base Methods: Full Fine-Tuning (FULL), Fixed Backbone (FIXED); 2) Domain Adaptation: Multi-Domain Adversarial Network (MDAN) (Pei et al., 2018), Moment Matching (M3SDA) (Peng et al., 2019), Bayesian Gaussian Mixture (GMDI) (Ling et al., 2024); 3) Single-Domain LLM Fine-tuning: LoRA (Hu et al., 2021), Adapter (Houlsby et al., 2019), LLaMA-Adapter (Zhang et al., 2024b), Q-LoRA (Dettmers et al., 2023), Tag-LLM (Shen et al., 2024); 4) Data Selection: INSTRUCTMINING (IT) (Cao et al., 2024), S2L (Yang et al., 2024b). Please refer to A.2 for more details on the baselines.
@@ 5.1 Experimental Results
Memory Usage. Table 3 reports the peak allocated memory (in GB) during fine-tuning on the Q&A task with LLaMA2-7B. We measure memory usage using a single NVIDIA A100 GPU. The FULL method requires the largest memory footprint due to updating all model parameters. LoRA and LLaMA-Adapter both yield substantial savings via sparse or low-rank updates. Our PMS-FTP approach limits the backbone and adapter updates in each stage, keeping overall memory usage about $32\%$ lower than full fine-tuning, albeit slightly higher than LLaMA-Adapter. Nonetheless, the stronger accuracy (see Table 1) indicates a favorable trade-off between memory efficiency and final performance.
@@ 6 Conclusion
In future work, we aim to adapt this stage-wise procedure to a continual learning setting, where new domains arrive sequentially, thereby offering a more flexible lifelong learning framework for large language models. Furthermore, combining the proposed method with pruning strategies (Lu et al., 2024b; Zhou et al., 2025; Li et al., 2025a) represents an interesting direction, especially when dealing with models that have a massive number of parameters.
@@ 10. Guidelines:
    • The authors are encouraged to create a separate “Limitations” section in their paper.
@@ 35. Guidelines:
    • The authors should answer “Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
@@ 60. Guidelines:
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
@@ A.2 Baselines
    • LLaMA-Adapter [Zhang et al., 2024b]: zero-init residual adapters.
@@ A.3 Implementation Details
where $V_{i}$ is the vocabulary set of $\mathcal{D}_{i}$, $\operatorname{Jacc}(V_{i},V_{j})=|V_{i}\cap V_{j}|/|V_{i}\cup V_{j}|$, and $\mu_{i}$ is the mean Sentence-BERT embedding of domain $i$. Both terms are in $[0,1]$; the average is therefore in $[0,1]$.
@@ B.7 Inference footprint after LoRA merging
After each stage we merge the finished LoRA into the frozen backbone, so only one 4-bit adapter is carried at inference. Table 15 shows equal-or-lower memory than a single-adapter baseline. Together with the accuracy gains in Table 1, merging achieves a strictly better accuracy-memory trade-off than training/keeping multiple adapters.
@@ B.10 Incremental addition of a new domain
After training on the original four domains, we add Legal-QA as a new stage and freeze prior adapters. Table 19 shows negligible forgetting on old domains and a gain over single-domain LoRA on the new domain. Disjoint, frozen adapters make PMS-FTP naturally amenable to one-shot domain extension without replay.
Table 19: One-shot incremental stage (add Legal-QA).